OpenStudy (anonymous):

The volume of a cone varies jointly as the square of its radius and its height. (V=k*r^2*h) If the volume of a cone is 9pi inches when the radius is 3 inches and its height is 3 inches, find the volume of a cone when the radius is 5 inches and the height is 9 inches. Express the answer in terms of pi. Given answer is 75 pi. How do I work this? Please help? Anyone?

OpenStudy (hoblos):

first you need to calculate k you have V=k*r^2*h so 9pi = k*3^2*3 so what's k ?

OpenStudy (anonymous):

that's where I am having the problem.

OpenStudy (anonymous):

If I could figure out the k, I'd be halfway to figuring this out. The other half is trying to figure out how to express the answer in terms of pi

OpenStudy (anonymous):

because I have another question exactly like it, and i feel like I am overlooking something incredibly simple

OpenStudy (hoblos):

\[9 \pi = k*3^2*3\] \[k = \frac{ 9 \pi }{ 3^2*3 }\]

OpenStudy (hoblos):

to express the answer in terms of pi, you have to keep k in terms of pi

OpenStudy (anonymous):

how do I do that?

OpenStudy (hoblos):

you just keep the pi and simplify the rest so \[k = \frac{ 9 }{ 9*3 } \pi\]

OpenStudy (hoblos):

can you simplify it ?

OpenStudy (anonymous):

k = 1/3 π

OpenStudy (hoblos):

correct

OpenStudy (anonymous):

hmmm....

OpenStudy (hoblos):

you didn't understand ?

OpenStudy (anonymous):

actually am trying to figure out how to plug the same value into another equation. want to see if the k i got from this answer will also be the same in the other question

OpenStudy (hoblos):

actually for a cone k is always 1/3 pi but that's only for a cone you can use this k to find the volume of the other cone mentioned in this question "radius is 5 inches and the height is 9 inches"

OpenStudy (anonymous):

you're a life saver! i wish I could give you 2 medals for this question.

OpenStudy (anonymous):

since the other question is also about a cone, then yes, I can do the math and move on. i simply could not figure out how to manipulate the math to make pi behave

OpenStudy (hoblos):

so you are able to do it now ?

OpenStudy (anonymous):

let me see. V=1/3 *r^2*h my next question is ?π=1/3*4^2*3

OpenStudy (hoblos):



OpenStudy (anonymous):

1/3π*4^2*3 1/3π*16*3 wouldn't the 1/3 and the 3 cross cancel each other out and leave me with 16π?

OpenStudy (hoblos):

yeah correct

OpenStudy (anonymous):

thankee-sai!

OpenStudy (hoblos):

glad to help! :)

OpenStudy (anonymous):

now let's see if I can manage to pull off a somewhat passing grade on this test :/

OpenStudy (hoblos):

i hope you can